Database last updated on Thu, 05 Feb 2026 22:10:53 +0000 by localhost

Critical warnings from b-files' comments for days between 2022174 and 2022174 (0 sec, 11 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 174 17 16 15 Motor #3: appears stuck with a reference sensor activated.31B17422.190TORON#19005:29:531-1
rl:2022 174 17 34 27 New value set for watermark LOOP.RUN.TIME: 653.999996 millisec.53B17422.190TORON#19005:30:011-1
rl:2022 174 17 34 27 New value set for watermark LOOP.RUN.TIME: 653.999996 millisec.53B17422.190TORON#19005:30:011-1
hp:Spectrometer alignment trouble.31B17422.190TORON#19011:23:081-1
hg:Lamp not detected f1= 1 f4= 031B17422.190TORON#19011:25:491-1
hg:Lamp not on... hg test terminated F1= 1 F4= 031B17422.190TORON#19011:25:491-1
hg:No intensity detected. Check PMT.11B17422.190TORON#19012:44:161-1
menu:Brewer failed to respond 5 times21B17422.190TORON#19012:59:431-1
si:Azimuth drive turned OFF in op12B17422.190TORON#19017:23:541-1
pd:Command file pd not found --- continuing11B17422.190TORON#19017:25:241-1
ds:Suppressed oscillation of filter wheel.175B17422.190TORON#19019:23:541-1

Instrument constant changes for the week before 2022174 (1 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
1902022174187.8

Non-critical warnings from b-files' comments for days between 2022174 and 2022174 (0 sec, 18 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
zs:ZS measurement terminated15B17422.190TORON#19000:20:111-1
rl:2022 174 17 16 14 Reset Tepid -> Tepid from 2022 174 17 11 40 until 2022 174 17 16 09 . Counts: 1,013B17422.190TORON#19005:29:511-1
rl:2022 174 17 16 15 Motor #3: appears stuck with a reference sensor activated.13B17422.190TORON#19005:29:531-1
rl:2022 174 17 16 15 Motor #10: appears stuck with a reference sensor activated.13B17422.190TORON#19005:29:561-1
rl:2022 174 17 16 15 Motor #11: appears stuck with a reference sensor activated.13B17422.190TORON#19005:29:591-1
rl:2022 174 17 43 36 Reset Warm -> Warm from 2022 174 17 43 33 until 2022 174 17 43 33 . Counts: 0,113B17422.190TORON#19005:30:101-1
rl:2022 174 17 45 46 Bad command string. Command string buffer cleared.13B17422.190TORON#19005:30:121-1
rl:2022 174 17 45 46 Bad command string. Command string buffer cleared.13B17422.190TORON#19005:30:121-1
rl:2022 174 17 45 46 <13B17422.190TORON#19005:30:141-1
ds:DS intensity too low ( 1/cy), skipping.499B17422.190TORON#19009:32:241-1
ab_sk:HOME key pressed499B17422.190TORON#19011:13:561-1
dsp:Doing dispersion test with Cd (UV port)399B17422.190TORON#19011:21:351-1
hg:DEL key pressed199B17422.190TORON#19013:40:091-1
checkhg:Last HG had to adjust the micrometer by at least 1299B17422.190TORON#19014:41:121-1
si:User didn't confirm doing sighting. Quitting.15B17422.190TORON#19017:24:581-1
pd:Command file pd not found --- continuing199B17422.190TORON#19017:25:24continuing1-1
si:New SI, HC = -6, NC = 120 Old = -8 , 120299B17422.190TORON#19017:30:111-1
pd:Going to menu299B17422.190TORON#19017:37:261-1

Measurements and tests summary.


_jday_loc_serial_skcaodedsdtedfzhghprssluvzs
174TORONTO#190ecsc,menu,Next schedule will be calo3 (now calo3 ZS= 93 ZS= 94.015),Starting schedule calo3,Starting schedule calsc2829121342814176
(0 sec)

Relative and absolute humidity for day 2022174 (2 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
19037.48.1